Use the quotient rule to simplify the expression:
(-11)21
-------
(-11)19
Note- The problem reades -11 to the 21 power divide by -11 to the 19 power.
Thanks,
Pat
: Use the quotient rule to simplify the expression:
(-11)21
-------
(-11)19
Note- The problem reades -11 to the 21 power divide by -11 to the 19 power.
Thanks,
Pat

Answer by Fombitz(1740)

When you divide exponential expressions of the same base, you subtract the exponents.
x^M/x^N=x^(M-N)
In your case x=-11 and M and N are 21 and 19,
(-11)^21/(-11)^19=(-11)^(21-19)
(-11)^21/(-11)^19=(-11)^(2)
(-11)^21/(-11)^19=121